# A fully Bayesian approach for the imputation and analysis of derived outcome variables with missingness

### v1

@@ 3 Description of the approach
    For $m$ in 1,…,$M$, draw $S$ iid samples from the $(Z_{1:P}|X=0,\phi^{[m]})$ distribution: $Z^{*[m,s]}_{1:P}|_{X=0}$, for $s$ in 1,…, $S$; and $S$ iid samples from the $(Z_{1:P}|X=1,\phi^{[m]})$ distribution: $Z^{*[m,s]}_{1:P}|_{X=1}$, for $s$ in 1,…, $S$.
@@
    Apply the deterministic function $f$ to the samples obtained in step 2 to obtain $S$ iid samples from the $(Y|X=0,\phi^{[m]})$ distribution: $Y^{*[m,s]}|_{X=0}$, for $s$ in 1,…, $S$; and $S$ iid samples from the $(Y|X=1,\phi^{[m]})$ distribution: $Y^{*[m,s]}|_{X=1}$, for $s$ in 1,…, $S$.
@@
    Use the samples obtained in step 3 to numerically approximate a draw of $\theta^{[m]}$, for $m$ in 1,…,$M$. Note that $\theta^{[m]}\approx h(\phi^{[m]})$ approaching equality with sufficiently large $S$. For instance, in our earlier example we had $\theta=\textrm{E}(Y|X=1)-\textrm{E}(Y|X=0)$, and would substitute the two expectations with their respective sample means, such that:
    ${\theta}^{[m]}=\frac{\sum_{s=1}^{S}{Y^{*[m,s]}|_{X=1}}}{S}-\frac{\sum_{s=1}^{S}{Y^{*[m,s]}|_{X=0}}}{S},$ (3)
    for $m$ in 1,…,$M$.
Having obtained $M$ posterior samples of $\theta$ following these four steps, one can then calculate a posterior median/mean and credible interval for $\theta$, as is standard practice in a Bayesian analysis. Both $M$ and $S$ should be sufficiently large that both sources of Monte Carlo error are negligible. In the next section we demonstrate how this approach can be applied in the motivating application.
4 Application to data: Estimating the risk of microcephaly
@@
Zika virus infection (ZIKV) during pregnancy is known to be associated with an increased risk of fetal congenital malformations, including microcephaly, a birth condition in which a baby’s head is smaller than expected when compared to babies of the same sex and age. Formally, microcephaly is defined as a measure of head circumference at birth of more than 2 standard deviations below average for sex and gestational age; see Harville et al. (2020). (Average values for head circumference at birth can be determined by referencing the standards calculated by the intergrowth-21st research network; see Villar et al. (2014)). Therefore, in a healthy population, we would anticipate 2.28% of newborns being categorized as having microcephaly (assuming head circumference in the general population is Gaussian).
In order to determine the risk of microcephaly associated with ZIKV, observational data has been collected by numerous studies from infected pregnant women (Zika Virus Individual Participant Data Consortium 2020). These data typically suffer from substantial missingness.
The outcome of interest, microcephaly ($Y$), is a derived binary variable, based on three source variables: (1) sex ($Z_{1}$=0 for ‘male’ and $Z_{1}$=1 for ‘female’), (2) gestational age ($Z_{2}$, measured in weeks), and (3) head circumference ($Z_{3}$, measured in centimetres). Each of these three source variables is subject to missingness. The exposure variable is maternal ZIKV status ($X$) but in this example the data we consider only include subjects for whom $X=1$. Specifically, we consider a simple artificial dataset based on what is typically observed in observational studies. The data is simulated with a “true” risk of microcephaly of $\theta$ = 11.69%. Missing $Z_{1}$ values are attributed completely at random, missing $Z_{2}$ values are more likely for those with very small values of $Z_{3}$, and missing $Z_{3}$ values are more likely for those with very large values of $Z_{2}$.
@@
The data consist of $Z_{1}$, $Z_{2}$, and $Z_{3}$ values (when available) from $N=1,800$ newborns; see Figure 2. There are 781 individuals who have at least one value missing; see Venn diagram in Figure 3. Microcephaly status, $Y$, for a newborn can be calculated by first deriving their “z-score” based on their head circumference, gestational age and sex. (The “igb_hcircm2zscore” function in the growthstandards R package can do this easily using the intergrowth standards). If a newborn’s z-score is below $-2$, they are classified as having microcephaly. As such, while the $f()$ function is known and can be easily applied to the source variables to obtain the binary derived variable, it is not easily written out.
@@
$Y=1|X=1\sim\textrm{Bernoulli}(\theta).$ (4)
We set a $\textrm{Beta}(1.26,2.39)$ prior for $\theta$ (see top panel of Figure 4), corresponding to an a priori belief that $\textrm{Pr}(\theta<31.5\%)=50\%$.
We fit the Bernoulli model to the complete case data ($N=1,019$) and obtained an estimate of: $\hat{\theta}=9.49\%$, 95%CrI= [7.84%, 11.39%] (posterior median and equal-tailed 95% credible interval). We also fit the Bernoulli model to the entire dataset ($N$= 1,800) and followed Gelman’s three-step procedure with SVL imputation. Specifically, we used MICE specifying the “logreg” method for $Z_{1}$ and the “normal” method for $Z_{2}$ and $Z_{3}$. For each of 50 imputed datasets, we fit the Bernoulli model using MCMC with $M$ = 5,000 draws and after combining all 250,000 draws together, we obtained an estimate of: $\hat{\theta}=12.73\%$, 95%CrI= [11.06%, 14.56%].
@@
Finally, head circumference ($Z_{3}$) is known to be approximately Gaussian and increases in an approximately quadratic way with gestational age (see Figure 2C in Villar et al. (2014), and Figure 8 in the Appendix). In a ZIKV infected population, infection will cause a certain proportion of infants to have smaller heads. As in Kalmin et al. (2019), we define a Normal mixture distribution for $Z_{3}$:
@@
$\mu$ $\sim\textrm{Normal}(0,0.1),\quad\beta_{0}\sim\textrm{Normal}(33.912,0.1),$
$\beta_{1}$ $\sim\textrm{Normal}(-0.450,0.1),\quad\beta_{2}\sim\textrm{Normal}(0.399,0.1),$
and $\quad\beta_{3}\sim\textrm{Normal}(-0.016,0.1).$
@@
Finally, for the two parameters that define the impact of ZIKV, we set the following priors:
@@
The truncated Normal prior for $\kappa$ was chosen in an effort to help address issues with identifiability that are common when fitting Bayesian mixture models; see Betancourt (2017).
These priors correspond to an a priori belief that $\textrm{Pr}(\textrm{E}(Y|X=1)<31.5\%)=0.5$. The bottom panel of Figure 4 plots the implied prior distribution for $(\textrm{E}(Y|X=1)$ (i.e., for the risk of microcephaly) and Figure 5 plots the implied distribution of the z-score from 12 random draws of the prior.
Fitting the BsNmN model to the complete case data ($N=1,019$) we obtain an estimate of: $\hat{\theta}=9.20\%$, 95%CrI= [7.62%, 10.94%]. We also applied the proposed approach to fit the BsNmN model to the entire dataset ($N=1,800$) as detailed in Section 2 with $S$=5,000 and obtained an estimate of: $\hat{\theta}=11.72\%$, 95%CrI= [10.18%, 13.30%]. Figure 6 plots the implied distribution of the z-score from the prior and posterior. Figures 9 and 10 in the Appendix show the trace plots of the MCMC.
Figure 4: Lower panel shows histogram of a sample from the “implied prior” on $\textrm{E}(Y|X=1)$ used in the BsNmN model. Upper panel shows histogram of a sample from the Beta(1.33, 2.42) prior placed on $\theta$ used in the Bernoulli model. The numbers 1.33 and 2.42 were specifically chosen so that these two priors would be similar.
Figure 5: The implied distribution of the z-score from 12 random draws from the prior of the BsNmN model. For reference, the red curve corresponds to the density of the standard normal.
Data Model Estimate
Complete case data ($N=1,019$) Bernoulli $\hat{\theta}=9.49\%$, 95%CrI= [7.84%, 11.39%]
BsNmN $\hat{\theta}=9.20\%$, 95%CrI= [7.62%, 10.94%]
Entire dataset ($N=1,800$) Bernoulli $\hat{\theta}=12.73\%$, 95%CrI= [11.06%, 14.56%]
BsNmN $\hat{\theta}=11.72\%$, 95%CrI= [10.18%, 13.30%]
Table 4: Results from the microcephaly analyses (posterior medians and equal-tailed 95% credible intervals).
@@
Figure 7: The implied distribution of the $Z_{2}$ (gestational age) values from 50 imputation models (MICE with SVL imputation). For reference, the blue curve corresponds to the distribution of $Z_{2}$ values before any missigness was applied.

### v2

@@ 3 Description of the approach
    For $m$ in 1,…,$M$:
    1. (a)
      For $d$ in 1,…,$D$:
      1. i.
        If the $d$-th target population is defined in terms of unknown parameters, $\gamma^{d}$, these must be estimated from the data. Obtain a draw from $\gamma^{d}|(\phi_{source},W_{1,expl}^{\{mis\}})^{[m]}$: $\gamma^{d[m]}$.
      2. ii.
        Draw $S$ (where $S$ is a large number) iid samples from the $(W_{source}|\gamma^{d[m]})$: $W^{*d[m,s]}_{source}$, for $s$ in 1,…, $S$. Note: $\dim(W^{*d[m,s]}_{source})=(1,P)$.
      3. iii.
        For $s$ in 1,…,$S$: Apply the deterministic function $f()$ to $W^{*d[m,s]}_{source}$ to obtain a sample from the $(Y|\gamma^{d[m]})$ distribution: $Y^{*d[m,s]}=f(W^{*d[m,s]}_{source})$. Note: $\dim(Y^{*d[m,s]})=(1,1)$.
      4. iv.
        Calculating (approximately) the sample mean:
        $\textrm{E}^{*}_{d}\{\textrm{E}(Y|W_{expl})\}^{[m]}\approx\frac{1}{S}\sum_{s=1}^{S}{Y^{*d[m,s]}}$.
    2. (b)
      Obtain ${\theta}^{[m]}$ by combining $\textrm{E}^{*}_{1}\{\textrm{E}(Y|W_{expl})\},...,\textrm{E}^{*}_{D}\{\textrm{E}(Y|W_{expl})\}$ according to the definition of the estimand of interest.
Having obtained $M$ posterior samples of $\theta$ following the four steps, one can calculate the posterior median/mean and credible interval for $\theta$, as is standard practice in a Bayesian analysis. Note that both $M$ and $S$ should be sufficiently large that both sources of Monte Carlo error are negligible.
In the simple illustrative example, the $D=2$ target populations were defined by $X=0$ and $X=1$. Therefore the $m$-th posterior sample of $\theta$ is approximated as:
$\theta^{[m]}=\frac{1}{S}\sum_{s=1}^{S}{Y^{*2[m,s]}}-\frac{1}{S}\sum_{s=1}^{S}{Y^{*1[m,s]}},$ (4)
where $Y^{*1[m,s]}$ is a sample from the distribution of $(Y|\phi_{source}^{[m]},X=0)$ and $Y^{*2[m,s]}$ is a sample from the distribution of $(Y|\phi_{source}^{[m]},X=1)$.
4 Applied examples
4.1 Multiple linear regression of the body mass index of Dutch boys
Background
The “Growth of Dutch boys” dataset is available as an example dataset within the ‘mice’ package for R (van Buuren et al. 2015) and is often used to demonstrate how various imputation methods can be applied. Here we use this dataset to illustrate how the approach outlined in Section 3 could be applied for determining the difference in logBMI between boys from the “city” and boys from outside of the “city”, when controlling for age.
The outcome of interest, $Y=\textrm{log}(\textrm{BMI})$, is a derived binary variable, based on two source variables: (1) height ($Z_{1}$, measured in centimeters), (2) weight ($Z_{2}$, measured in kilograms). This is derived as follows:
$Y=\textrm{logBMI}=f(Z_{1},Z_{2})=\textrm{log}\Big{(}\frac{Z_{2}}{({Z_{1}/100})^{2}}\Big{)}.$ (5)
The exposure variable is regional status ($X$, equal to 1, if the boy is from the “city” and equal to 0 otherwise). We also consider the variable age ($A$, measured in years) as highly predictive of $Z_{1}$, $Z_{2}$, and $Y$.
The data
The data consists of measurements on a representative sample of $n=537$ boys between the ages of 1 and 18 years (inclusive), of which approximately 9% are from the “city” (we restricted the original “Growth of Dutch boys” dataset dataset ($n=748$) to the 1–18 age range to allow for reasonable modeling of how log(BMI) changes with age). Figure 1 plots the data among the complete cases. All variables, except age ($A$), are subject to a small amount of missingness. The exposure variable, “city” ($X$), is missing for 1 subject; height is missing for 18 subjects; and weight is missing for 2 subjects; see Venn diagram in Figure 2. Overall, 19 out of the 537 of observations (3.5%) have at least one missing value. Following the notation defined in Section 3, we have that $W_{1}=\{Z_{1},Z_{2},X\}$, $W_{2}=\{A\}$, $p=2$, $W_{source}=\{Z_{1},Z_{2}\}$, $q=2$, and $W_{expl}=\{X,A\}$.
The primary objective of the study is to estimate the difference in mean logBMI between boys from inside the “city” and boys from outside of the “city”, when controlling for age. A secondary objective is to make probabilistic predictions of logBMI for boys given their age and regional status.
Figure 1: The complete case data for the Dutch boys analysis ($n$=518) with violet dots corresponding to boys from outside the city ($X=0$) and green dots corresponding to boys from the city ($X=1)$. Lines drawn from least-squares estimates.
Figure 2: Venn diagram for missingness in Dutch boys dataset ($n$=537).
The estimand
Based on the primary objective, the estimand of interest is the difference in mean logBMI between boys from the “city” and boys from outside of the “city”, when controlling for age, and this can be defined as a function of the conditional expectation of $Y$ given $D=2$ target populations:
[…]
where $f_{A}(a)$ corresponds to the distribution of ages among boys in the Netherlands. Since the data supposedly consist of a representative sample of Dutch boys, we assume that $f_{A}(a)$ also coincides with the sample distribution of ages.
Two models
The univariate normal model involves considering the logBMI of each subject as the outcome of a simple univariate Normal model:
$(Y|X=x,A=a)\sim\textrm{Normal}(\beta_{0}+\beta_{1}x+\beta_{2}a+\beta_{3}xa+\beta_{4}a^{2},\sigma^{2}).$ (7)
This model assumes a quadratic relationship between BMI and age, a reasonable approximation based on published growth curves (Kuczmarski 2000) and allows for the possibility that age is an effect modifier with respect to the effect of “city”. For the univariate normal model, we set standard Normal priors, $\textrm{Normal}(0,1)$, for the regression coefficient parameters $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$ and $\beta_{4}$, an exponential prior, $\textrm{Exponential}(1)$, for the standard deviation parameter, $\sigma$, and a uniform prior, $\textrm{Uniform}[0,1]$, for $\pi$.
The bivariate normal model considers height and weight as two correlated outcomes such that:
$(\textrm{log}(Z_{1}),\textrm{log}(Z_{2})|X=x,A=a)$ $\sim\text{Normal}(\boldsymbol{\mu_{i}},\boldsymbol{\Sigma}),$ (8)
where
$\boldsymbol{\mu_{i}}$ $=\begin{pmatrix}\alpha_{0}+\alpha_{1}r+\alpha_{2}a+\alpha_{3}xa+\alpha_{4}a^{2}\\ \gamma_{0}+\gamma_{1}x+\gamma_{2}a+\gamma_{3}xa+\gamma_{4}a^{2}\end{pmatrix}\textrm{and}\quad\quad$ (9)
$\boldsymbol{\Sigma}$ $=\begin{pmatrix}\tau_{Z1}^{2}&\rho\tau_{Z1}\tau_{Z2}\\ \rho\tau_{Z1}\tau_{Z2}&\tau_{Z2}^{2}\\ \end{pmatrix}.$
and:
$R\sim\textrm{Bernoulli}(\pi).$ (10)
We set standard Normal priors, $\textrm{Normal}(0,1)$, for all the regression coefficient parameters $\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$, $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$; exponential priors, $\textrm{Exponential}(1)$, for the standard deviation parameters, $\tau_{Z1}$ and $\tau_{Z2}$, a uniform prior, $\textrm{Uniform}[-1,1]$, for the correlation parameter, $\rho$, and another uniform prior, $\textrm{Uniform}[0,1]$, for $\pi$.
We can define the estimand of interest analytically as a function of the parameters of the univariate normal model:
[…]
$=\beta_{1}+\beta_{3}\textrm{E}(A),$ (12)
where $\phi_{univ}=\{\beta_{0},\beta_{1},\beta_{2},\beta_{3},,\beta_{4},\sigma\}$. Since the data is assumed to be a representative sample from the target population, and since we are choosing not to model the distribution of age (not required because there are no missing values for age), we can substitute $\textrm{E}(A)$ for the sample mean, $\bar{A}$.
Coincidentally, we can define the estimand as a function of the parameters of the bivariate normal model (see details in the Appendix):
[…]
$=\gamma_{1}-2\alpha_{1}+(\gamma_{3}-2\alpha_{3})\textrm{E}(A),$ (13)
where $\phi_{biv}=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\tau_{Z1},\tau_{Z2},\rho\}$, and we once again can substitute $\textrm{E}(A)=\bar{A}$.
Methods and Results
We conducted the analysis with four different approaches:
  1. 1.
    Fitting the univariate normal model to the complete case data and applying the $h_{1}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$).
  2. 2.
    Fitting the univariate normal model with Gelman’s three-step approach with “on-the-fly” multiple imputation (including a quadratic term for age and age-city interaction term in the imputation model), and applying the $h_{1}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$, and $K=50$ imputed datasets).
@@ 4.1 Multiple linear regression of the body mass index of Dutch boys
    Fitting the bivariate normal model and applying $h_{2}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$).
@@
    Fitting the bivariate normal model and applying the proposed method detailed in Section 3 (with $M=2,000$ , and $S=2,000$).
Both the univariate model and bivariate model were fit using JAGS with 1,000 burn-ins (Plummer 2004). R and JAGS code is available in the Appendix.
Results are listed in Figure 3. The estimate from the complete case analysis is similar to the estimates obtained when using the the bivariate model with a wider credible interval, suggesting that one can gain precision by incorporating the small number of observations with missing values into the analysis. However, one obtains a wider credible interval when using the univariate normal model with Gelman’s three-step approach with “on-the-fly” multiple imputation, suggesting that precision can also be lost. Finally, the estimate obtained with fitting the bivariate normal model and applying $h_{2}()$ function is slightly more precise than the estimate obtained with fitting the bivariate normal model and applying the proposed method, suggesting that there is a small cost to the numerical approximation (even with the large $S=2,000$). With respect to computational time, the proposed method is substantially more time consuming than the other methods.
Figure 3: Results for the analyses of the Dutch boys dataset. Estimates list the posterior median estimates (i.e., the median of the posterior estimates of the mean difference in logBMI) and equal-tailed posterior 95%CrIs. Labels indicate computational time required for each method in minutes.
4.2 Estimating the risk of microcephaly
@@
Zika virus infection (ZIKV) during pregnancy is known to be associated with an increased risk of fetal congenital malformations, including microcephaly, a birth condition in which a baby’s head is smaller than expected when compared to babies of the same sex and age. Formally, microcephaly is defined as a measure of head circumference at birth of more than 2 standard deviations below average for sex and gestational age; see Harville et al. (2020). (Average values for head circumference at birth are determined by referencing the standards calculated by the intergrowth-21st research network; see Villar et al. (2014)). Therefore, in a healthy population, we would anticipate 2.28% of newborns being categorized as having microcephaly (assuming head circumference in the general population is Gaussian).
In order to determine the risk of microcephaly associated with ZIKV in mothers, observational data has been collected by numerous studies from infected pregnant women (Zika Virus Individual Participant Data Consortium 2020). These data typically suffer from substantial missingness.
The outcome of interest, microcephaly ($Y$), is a derived binary variable, based on three source variables: (1) sex ($Z_{1}$=0 for ‘male’ and $Z_{1}$=1 for ‘female’), (2) gestational age ($Z_{2}$, measured in weeks), and (3) head circumference ($Z_{3}$, measured in centimetres). Each of these three source variables is subject to missingness. The estimand of interest is the probability of microcephaly ($\theta=\textrm{E}(Y)$) given then assumed $D=1$ target population from which the data is sampled.
@@
We consider a simple synthetic dataset based on what is typically observed in observational studies. The purpose of this analysis is to demonstrate all of the steps required when following each of the different approaches for dealing with missingness based on the problem that initially motivated our research. As such, we stress that the specific details of how the data were created are less important and acknowledge that a different dataset would no doubt lead to different results. In brief, the data is simulated with a “true” risk of microcephaly of $\theta$ = 11.69%. The data represent $Z_{1},Z_{2},Z_{3}$ measurements on $n=1,800$ newborns with at least one of $Z_{1},Z_{2},Z_{3}$ observed; Figure 4 shows $Z_{3}$ and $Z_{2}$ among those with both observed. Missing $Z_{1}$ values are attributed completely at random, missing $Z_{2}$ values are more likely for those with very small values of $Z_{3}$, and missing $Z_{3}$ values are more likely for those with very large values of $Z_{2}$.
There are 781 individuals who have at least one value missing; see Venn diagram in Figure 5. Microcephaly status, $Y$, for a newborn can be calculated by first deriving their “z-score” based on their head circumference, gestational age and sex. (The “igb_hcircm2zscore” function in the growthstandards R package can do this easily using the intergrowth standards). If a newborn’s z-score is below $-2$, they are classified as having microcephaly. As such, while the $f()$ function is known and can be easily applied to the source variables to obtain the binary derived variable, it is not easily written out.
@@
$Y\sim\textrm{Bernoulli}(\theta).$ (14)
We set a $\textrm{Beta}(1.26,2.32)$ prior for $\theta$ (see top panel of Figure 6), corresponding to an a priori belief that $\textrm{Pr}(\theta<31.5\%)=50\%$. (This prior for was selected so that the two models have (implied) priors matching as closely as possible; see Figure 6).
We fit the Bernoulli model to the complete case data ($n=1,019$) and obtained an estimate of: $\hat{\theta}=9.18\%$, 95%CrI= [7.60%, 10.92%] (posterior median and equal-tailed 95% credible interval). We also fit the Bernoulli model to the entire dataset ($n$= 1,800) and followed Gelman’s three-step procedure with SVL imputation. Specifically, we used MICE specifying the “logreg” method for $Z_{1}$ and the “normal” method for $Z_{2}$ and $Z_{3}$. For each of $K=50$ imputed datasets, we fit the Bernoulli model using MCMC with $M$ = 5,000 draws and after combining all 250,000 draws together, we obtained a posterior median estimate of: $\hat{\theta}=12.70\%$, with equal-tailed 95%CrI= [11.00%, 14.51%].
@@
Finally, head circumference ($Z_{3}$) is known to be approximately Gaussian and mean head circumference increases in an approximately quadratic way with gestational age (see Figure 2C in Villar et al. (2014), and Figure 11 in the Appendix). In a ZIKV infected population, infection is thought to cause a certain proportion of infants to have smaller heads. As in Kalmin et al. (2019), we define a Normal mixture distribution for $Z_{3}$:
@@
$\mu\sim\textrm{Normal}(0,0.1),$
$\beta_{0}\sim\textrm{Normal}(33.912,0.1),\quad$ $\beta_{1}\sim\textrm{Normal}(-0.450,0.1),$
$\beta_{2}\sim\textrm{Normal}(0.399,0.1),\quad\textrm{and}$ $\beta_{3}\sim\textrm{Normal}(-0.016,0.1).$
@@
Finally, a truncated Normal prior for $\kappa$ was chosen in an effort to help address issues with identifiability that are common when fitting Bayesian mixture models (Betancourt 2017), and a uniform prior was chosen for $w$:
@@
These priors correspond to an a priori belief that $\textrm{Pr}(\textrm{E}(Y)<31.5\%)=0.5$. The bottom panel of Figure 6 plots the implied prior distribution for $\textrm{E}(Y)$ (i.e., for the risk of microcephaly) and Figure 7 plots the implied distribution of the z-score from 12 random draws of the prior.
Fitting the BsNmN model to the complete case data ($n=1,019$) we obtain an estimate of: $\hat{\theta}=9.18\%$, 95%CrI= [7.60%, 10.92%]. We also applied the proposed approach to fit the BsNmN model to the entire dataset ($n=1,800$) as detailed in Section 2 with $S$=5,000 and obtained an estimate of: $\hat{\theta}=11.68\%$, 95%CrI= [10.16%, 13.36%]. Figure 9 plots the implied distribution of the z-score from the prior and posterior. Figures 12 and 13 in the Appendix show the trace plots of the MCMC.
Figure 6: Lower panel shows histogram of a sample from the “implied prior” on $\textrm{E}(Y)$ used in the BsNmN model for the microcephaly example analysis. Upper panel shows histogram of a sample from the Beta(1.26, 2.32) prior placed on $\theta$ used in the Bernoulli model. The numbers 1.26 and 2.32 were specifically chosen so that these two priors would be similar.
Figure 7: The implied distribution of the z-score from 12 random draws from the prior of the BsNmN model for the microcephaly example analysis. For reference, the red curve corresponds to the density of the standard normal.
Figure 8: Results from the microcephaly analyses (posterior medians and equal-tailed 95% credible intervals).